The probability that a randomly selected 3-year-old male feral cat
will live to be 4 years old is 0.99612.
(a) What is the probability that two randomly selected 3-year-old male feral cats will live to be 4-years old?
(b) What is the probability that five randomly selected 3-year-old male feral cats will live to be 4-years old?

Answer by mathmate(429)

Question:
The probability that a randomly selected 3-year-old male feral cat
will live to be 4 years old is 0.99612.
(a) What is the probability that two randomly selected 3-year-old male feral cats will live to be 4-years old?
(b) What is the probability that five randomly selected 3-year-old male feral cats will live to be 4-years old?

Solution:
Use the multiplication rule. The probability of successes of all n independent experiments each with probability p equals p^n.
(a) p=.99612, n=2, so P(2)=0.99612^2=0.99226
(b) p=.99612, n=5, so P(5)=0.99612^5=0.98075